OpenStudy (anonymous):

Antonio has budgeted an average of $45 a month for entertainment. For the first five months of the year he has spent $48, $39, $60, $48, and $33. How much can Antonio spend in the sixth month without exceeding his average budget?

OpenStudy (amistre64):

sum(48,39,60,48,33) + x ----------------------- = 45 6 lets call the sum S S + x ----- = 45 6 S + x = 45(6) x = 45(6) - S

OpenStudy (amistre64):

and if i did it right; I get $42 ?

OpenStudy (anonymous):

hmm its wronq :/

OpenStudy (amistre64):

dbl check: (48+39+60+48+33+x)/6 = 45 (48+39+60+48+33+42)/6 = 45 270/6 = 45 45 = 45

OpenStudy (amistre64):

45*6 = 270 270 - (48+39+60+48+33) = 42 no matter how I do it; its always 42

OpenStudy (amistre64):

another way is to see what he has spent like this: $48 = 3 over $39 = 6 under $60 = 15 over $48 = 3 over $33 = 12 under 3 - 6 = -3 +15 = 12 +3 = 15 -12 = 3 hes 3 bucks over at the moment; 45-3 = 42 dollars to get back on track